i need complete solution for this logical proof
AB+CD=(A+C)(A+D)(B+C)(B+D)
Where:
A+BC=(A+B)(A+C)
A+AB=A
A+B=B+A
A= AA
A+A=A

Answer by robertb(5830)

(A+C)(A+D)(B+C)(B+D)
=((A+C)(A+D))((B+C)(B+D)) ASSUMING associativity
= (A + CD)(B + CD) from A+BC=(A+B)(A+C)
= (CD + A)(CD + B) from A+B=B+A
= CD + AB from A+BC=(A+B)(A+C)